OpenStudy (anonymous):

find the domain of the fuction and express the answer in interval notation: g(x)=10/(x+7)

OpenStudy (anonymous):

i believe the answer to this is (-7,\[\infty\])

OpenStudy (anonymous):

no, the only number you are not permitted is what would make the denominator 0

OpenStudy (anonymous):

so you were right to single out -7, but that is the only number not permitted. x could be less than -7 as well as greater

OpenStudy (anonymous):

i.e. all real numbers except -7 is your answer

jimthompson5910 (jim_thompson5910):

Domain is \[\large \left(-\infty,-7\right)\cup\left(-7,\infty\right)\] Basically, it's the set of all real numbers BUT x can't equal -7